OpenStudy (anonymous):

How long would it take $3500 to grow to $8400 if the annual rate is 4.9% and interest is compounded monthly?

OpenStudy (anonymous):

\[8400=3500(1+\frac{.049}{12})^{12t}\] solve for t

OpenStudy (anonymous):

any idea how to do this? you need logs

OpenStudy (anonymous):

when it comes to logs i am beyond lost ! to simplify i am going to do whats in the () first

OpenStudy (anonymous):

ok the first thing you need handy is a calculator

OpenStudy (anonymous):

and the next thing before messing around in the parentheses is to divide both sides by 3500

OpenStudy (anonymous):

2.4

OpenStudy (anonymous):

i.e. start with \[2.4=(1+\frac{.049}{12})^{12t}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now with careful use of a calculator we can solve for \(12t\) in one step

OpenStudy (anonymous):

the idea is this to solve \[b^x=A\] you get \[x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

in this example \(A=2.4,b=(1+\frac{.049}{12})\)

OpenStudy (anonymous):

so you get \[12t=\frac{\ln(2.4)}{\ln(1+\frac{.049}{12})}\]

OpenStudy (anonymous):

of course we want \(t\) not \(12t\) so your actual answer is \[t=\frac{\ln(2.4)}{12\ln(1+\frac{.049}{12})}\]

OpenStudy (anonymous):

i get 17.9 carefully using parenthese and a calculator (ok wolfram) http://www.wolframalpha.com/input/?i=ln%282.4%29%2F%2812%28ln%281%2B.049%2F12%29%29%29

OpenStudy (anonymous):

to recap, the idea is to get it to look like \[b^x=A\]so you can write \[x=\frac{\ln(A)}{\ln(b)}\] the log of the total divided by the log of the base

OpenStudy (anonymous):

when you were putting the equation in to wolframaalpha.com why is there the 3 () at the end ?

OpenStudy (anonymous):

because ... let me go back and look

OpenStudy (anonymous):

ok i didn't really need it. here it is with few parentheses http://www.wolframalpha.com/input/?i=ln%282.4%29%2F%2812ln%281%2B.049%2F12%29%29

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

just be careful to put parenthese around the whole denominator and if you are using wolfram you can tell if it looks right because it shows you what it is computing yw

OpenStudy (anonymous):

yes that site is gonna be a huge help to me !